P(3a) = 5c3 ∙(0.61)3∙(0.39)2 = 0.3452... ~ 34.5%
False; the "or" is an additive property so the probability of rain or snow muse be greater than or equal to 0.65.
When rolling a pair of dice what is the probability that it will land on 11 or 12?
There are 36 possible combinations of rolling 2 die. To get an 11 or 12, the rolls would be 5, 6 or 6, 5 or 6,6 which is 3 possible correct outcomes, So, the probability of landing on an 11 or 12 is 3/36 or 1/12.
Function What is a graph sHow is a function if?
A graph has two axes, X and Y. A function can be seen on the graph based on the formula with X and Y representing certain properties in the formula.
What is theoretical probability of rolling a 2?
on a dice it would be 1/6, 0.166666666 (so on, unless rounded), and 16.667% (when rounded)
I will assume that you meant to ask what the probability is that exactly 25 of the computers Will require repair on a given day. This can be found by doing binompdf(150, .02, 25). This is a VERY small number (≈5.25*10^-16).
How many distinguishable words can be formed from the letters of the word casserole?
4-letter words
aces, acre, aero, alec, alee, ales, aloe, also, arco, arcs, ares, care, carl, cars, case, cees, cere, cero, cess, coal, cola, cole, cols, core, coss, earl, ears, ease, eels, else, eras, eros, eses, lace, lacs, lars, lase, lass, lear, leas, leer, lees, less, loca, lore, lose, loss, oars, ocas, olea, oles, oral, orca, orcs, ores, orle, osar, oses, ossa, race, rale, rase, real, recs, reel, rees, rocs, roes, role, rose, sacs, sale, sals, scar, seal, sear, seas, secs, seel, seer, sees, sels, sera, sere, sers, sloe, soar, sola, sole, sols, sora, sore
5-letter words
acres, alecs, aloes, arles, arose, arses, cares, carle, carls, carol, carse, cases, cease, ceorl, ceres, ceros, claro, class, clear, close, coals, colas, coles, coral, cores, corse, coses, crass, creel, cress, cross, earls, easel, eases, erase, erose, erses, escar, lacer, laces, laree, lares, laser, lases, lasso, lears, lease, leers, loess, lores, loser, loses, oases, ocrea, orals, orcas, orles, races, rales, rases, reals, reels, roles, roses, sales, saree, saros, scale, scare, scars, score, scree, seals, sears, seels, seers, serac, seral, seres, sloes, soars, socle, solar, soles, soras, sorel, sores
6-letter words
across, areole, caress, carles, carols, carses, ceases, ceorls, cereal, claros, clears, closer, closes, coaler, coarse, corals, corses, crases, crease, creels, creole, cresol, crosse, easels, erases, eroses, escars, lacers, larees, lasers, leaser, leases, lesser, lessor, losers, ocreae, oracle, rassle, reales, recess, recoal, relace, resale, reseal, resole, sarees, scaler, scales, scares, sclera, scores, screes, sealer, seracs, serosa, socles, solace, sorels
7-letter words
acerose, areoles, carless, caseose, cereals, claroes, classer, closers, coalers, creases, creoles, cresols, earless, escolar, lassoer, leasers, oarless, oracles, recoals, relaces, resales, rescale, reseals, resoles, scalers, sclerae, scleras, sealers, serosae, serosal, solacer, solaces
8-letter words
careless, coreless, escarole, escolars, lacrosse, rescales, sclerose, secalose, solacers
9-letter words
casserole, escaroles
A point is 30mm above the H P and 40mm in front of the V P draw its projection?
a point A is 40 mm above hp and 25 mm infront of vp. draw the frontview,topview and right side view
What is the probabilitty of flipping a coin three times and getting taills?
Each flip can come up in either one of 2 ways. So the number of different histories
of three flips is 2 x 2 x 2 = 8 .
One tails can come up in 3 ways in 3 flips: (H H T, H T H, or T H H).
Probability of one tails is 3/8 = 37.5% .
Two tails can happen in 3 ways in 3 flips: (H T T, T T H, or T H T).
Probability of two tails is 3/8 = 37.5% .
Three tails can only happen one way in 3 flips: T T T.
Probability of three tails = 1/8 = 12.5% .
No tails can only happen one way in 3 flips: H H H.
Probability of no tails = 1/8 = 12.5% .
The probability of any tails (1 or 2 or 3) is 7/8 = 87.5% .
The probability of 2 or 3 tails is 4/8 = 50% .
The probability of 1 or 3 tails is 4/8 = 50% .
The probability of no tails or any tails is 8/8 = 100% .
What does the probability of 0.618 mean?
It means that there is a 61.8% chance of the event occurring.
What is the probability that the sum of the dice is less than 6 and one die shows a 2?
Out of the 36 combinations, you have to count, or calculate, how many combinations are favorable. In this case:
1-2
2-1
2-2
2-3
3-2
That's 5/36. I am assumining that "one die shows a 2" means "at least one". Otherwise, if you mean "exactly one", you have to exclude the combination 2-2.
What is the probability of flipping a coin three times and it landing all three times?
100 percent. it will always land somewhere.
The theoretical probability that a tribute from 1,2,4 will win the Hunger Games is 25% chance. To find the experimental probability you would have to know which tributes won in the last 75 hunger games to know there chance of winning in the future.
Which of these events events would most likely result in the creation of large numbers of fossils?
Volcano
What is the probability of NOT rolling a sum of 12 with two dice?
The probabilityof ROLLING a sum of 12 is 1/36 ( 1 in 6 for each die = 1/6 x 1/6 = 1/36)
The probabilityof NOT ROLLING a sum of 12 is 35/36 ( 1 - 1/36)
What is the probability of getting a number less than four if you roll a die?
3/6 which equals 1/2.so 1/2 of the numbers less then 4 can be rolled on a die.
I think the equation is _X_X_X_X_=whatever, so it must be 3125.
How many ways can you put 4 different letters in order?
There are 24 ways (4 times 3 times 2 times 1) to put 4 different letters in order.
4x3x2 = 24
How do you explain why the probability that you will be the first person walk on the jmoojn in zero?
If "jmoojn" is the moon then the event has already happened and it was not you. So it is impossible and therefore the probability is 0.
What is the probability of rolling a sum of 1 explain?
It depends on "rolling a sum of 1" with what. One die, two dice or more?